# Open-Source Highly Parallel Electromagnetic Simulations for Superconducting Circuits

### v1

@@ I Introduction
The field of superconducting quantum circuits has emerged as one of the leading theoretical and experimental platforms in the pursuit of fault-tolerant quantum computing [1, 2]. Over the past two decades, this maturing field has experienced significant growth, highlighted by the evolving sophistication of experimental techniques and circuit design to improve qubit coherence times [3]. As the field continues to develop, focus is shifting from understanding the physics underlying superconducting circuit behaviour [4] to addressing engineering challenges, such as the design and optimization of device geometries [5]. Computational tools such as electromagnetic simulation software have become increasingly important for optimizing superconducting circuit design [6, 5]. Numerous frameworks have been proposed to simulate device behaviour and extract the circuit Hamiltonian using electromagnetic solvers. Among the most widely adopted are the lumped oscillator model (LOM) [7, 8], blackbox quantization (BBQ) [9], and the energy participation ratio (EPR) [10], which employ electrostatic, driven frequency domain, and eigenmode simulations, respectively. These approaches enable designers to evaluate, refine, and optimize circuit parameters prior to fabrication. However, performing such simulations remains challenging for researchers and engineers with limited access to proprietary electromagnetic solvers such as Ansys HFSS or COMSOL.
@@
  • Qiskit Metal [11] — an open-source framework developed by IBM for the design of superconducting quantum chips and devices.
  • Gmsh [12] — an open-source finite element mesh generator developed by Christophe Geuzaine and Jean-François Remacle for creating 3D meshes and models.
  • Palace (Parallel Large-scale Computational Electromagnetics) [13] — an open-source electromagnetic solver developed by Amazon Web Services (AWS) for HPC environments, enabling faster and more accurate large-scale simulations.
  • ParaView [14] — an open-source visualization and post-processing tool developed by Kitware for analyzing and rendering finite element simulation results.
@@ II Workflow
The operation of the SQDMetal API is shown in Fig. 1, providing a diagrammatic representation of our workflow. The process begins with Qiskit Metal, where the superconducting device is designed. Qiskit Metal features an extensive component library and enables users to programmatically develop circuit layouts in Python via Jupyter notebooks [11]. Qiskit Metal also supports community-driven development, allowing users to create and share custom components that can be integrated into new designs [15].
@@
After meshing, the SQDMetal API transfers the generated mesh file to Palace to perform finite-element electromagnetic simulations. Palace is developed on top of MFEM, an open-source high-performance C++ library for finite-element discretization, and is designed to operate across a wide range of platforms from laptops to supercomputers [16]. While proprietary software packages such as COMSOL Multiphysics and Ansys HFSS also support finite-element electromagnetic simulations on HPCs, there are few open-source alternatives. Palace addresses this gap by providing scalable, open-source computational electromagnetics capabilities. Within SQDMetal, we have integrated Palace’s functionality to perform the following types of simulations:
@@ III Mesh Convergence Study
where the last two designs were taken from the Qiskit Metal tutorial repository. All geometric and material parameters for each device were identical across solvers to ensure a fair comparison. The first two designs were defined on a $500\upmu\text{m}$-thick silicon substrate, while the third and fourth designs, used a $750\upmu\text{m}$-thick substrate. In all cases, the silicon substrate was assigned a relative permittivity of 11.45, a value commonly used for silicon at cryogenic temperatures [17]. The superconducting regions were modeled as zero-thickness surfaces with perfect electric conductor (PEC) boundary conditions. Each chip was enclosed in a rectangular box whose interior volume was defined as vacuum, with all bounding surfaces set to PEC to emulate placement within a metallic sample holder. The enclosing box was twice the height of the chip, while its width and length were 20% larger than the chip. The substrate was placed on the floor in the centre of the enclosing box, thereby grounding the bottom of the device. For the first two designs, the ends of the feedline were terminated with 50 $\Omega$ lumped ports to prevent signal reflections and ensure well-defined boundary conditions.
@@
Decreasing the size of mesh elements allows vector fields (e.g. the electric field in an eigenmode simulation) and scalar fields (e.g., the electric potential in an electrostatic simulation) to be captured with higher spatial resolution, resulting in more accurate solutions. For this benchmark study, we employ exclusively tetrahedral mesh elements. Within each element, the field quantities are interpolated using basis (shape) functions that approximate the field variation across the element [18, 19]. Higher-order basis functions improve the accuracy of field interpolation but increase computational cost (in both memory and time). The total number of global unknowns that an electromagnetic solver must compute are referred to as the degrees of freedom (DoF). The DoF directly determine the size of the global matrix system that is constructed at the beginning of the simulation and thus is an indicator of the computational resources required. Specifically, the size of the system matrices for eigenmode or electrostatic simulations is $n\times n$, where $n$ is the DoF. Increasing the number of mesh elements and order of the basis functions both increase the DoF.
@@
When implementing manual mesh refinement (MMR), the user increases the DoF by visually inspecting the electric field distribution and locally decreasing the mesh element size, that is, increasing the mesh element density in regions of strong electric field magnitude. SQDMetal provides specific commands that allow users to selectively refine regions of a device or individual circuit components. However, achieving convergence of a given parameter through MMR can be time-consuming and requires considerable user intuition. The alternative approach, adaptive mesh refinement (AMR), requires less user input while achieving excellent accuracy. In this method, the solver automatically refines the mesh by computing an error indicator that estimates the local deviation between the numerically computed electric field within an element and a smoothed field value calculated in post-processing [20, 21]. Elements exceeding a user-defined error threshold are subdivided into smaller elements until the estimated error falls below that threshold. AMR requires a reasonably well-refined initial seed mesh to be defined by the user in order to produce the most accurate results [5]. For both refinement schemes in our study, the polynomial order of the basis functions is held constant; therefore, the convergence behaviour observed in this study arises solely from changes in the mesh element size within regions of high electric field.
@@ III.1 Eigenmode Convergence Results
The results of the eigenmode mesh convergence study for each design are presented in Fig. 2. Each panel in the figure shows a schematic of the device geometry alongside its corresponding plot of resonant frequency as the DoF are increased. In the schematics, blue regions depict superconducting metals, modelled as a PEC, while gold regions indicate the exposed silicon substrate. The red rectangles between the capacitor pads are Josephson junctions modelled as lumped element ports. Each point in the plots represents a single eigenmode simulation conducted at a given DoF.
The data points in Fig. 2 correspond to results from different solvers and mesh refinement strategies. Results for Palace simulations using AMR and second order basis functions (AMR, O2) are shown as dark blue circles, while simulations conducted using Palace and MMR with fourth order basis functions (MMR, O4) are shown as light blue circles. Ansys HFSS simulations using AMR and second order basis functions (AMR, O2) are plotted in orange, while COMSOL Multiphysics simulations using MMR and second order basis functions (MMR, O2) are in red. For this study, we define a notional mesh element size as $m=\text{DoF}^{-1/3}$ [22] where $m\to 0$ as $\text{DoF}\to\infty$ to approach the continuum limit. The simulated resonant frequencies are plotted against this quantity to illustrate convergence of each solver toward a common frequency value. The corresponding total DoF for each simulation is indicated along the top axis of each plot for reference.
@@
The first benchmark design, shown in Fig. 2a, is a single coplanar waveguide (CPW) resonator. This geometry was selected as a baseline case, since a $\lambda/2$ resonator is a common component in superconducting circuits. The corresponding plot in Fig. 2b shows the convergence of the simulated resonant frequency. The resonator was designed to operate at 10 GHz based on the analytical expression for a half-wavelength CPW resonator [23]:
@@
where $f$ is the resonant frequency of the fundamental mode, $c$ is the speed of light in vacuum, $\ell$ is the length of the resonator and $\varepsilon_{\text{eff}}$ is the effective permittivity that accounts for field penetration into both vacuum and the substrate. The effective permittivity was approximated as $\varepsilon_{\text{eff}}=(\varepsilon_{\text{Si}}+\varepsilon_{\text{vac}})/2$, where $\varepsilon_{\text{Si}}=11.45$ [17] and $\varepsilon_{\text{vac}}=1$. Substituting $f=10$ GHz yields a resonator length of 6.012 mm.
@@
The second benchmark design, a coplanar waveguide (CPW) resonator capacitively coupled to a feedline, is shown in Fig. 2c, with corresponding eigenmode simulation results presented in Fig. 2d. This configuration introduces an additional degree of complexity relative to the single-resonator case via a capacitive coupling to a transmission feedline. For this $\lambda$/2 resonator, a length of 8.475 mm was used, which, according to Eq. (1), yields an expected fundamental frequency of approximately 7.089 GHz. As shown in Fig. 2d, at the highest DoF, Palace (MMR, O4) arrived at 7.095 GHz, and Palace (AMR, O2) produced 7.094 GHz. COMSOL Multiphysics and Ansys HFSS produced frequencies of 7.092 GHz (0.028%) and 7.101 GHz(0.127%), respectively.
@@
The next benchmark design, shown in Fig. 2e, consists of a transmon capacitively coupled to a coplanar waveguide (CPW) resonator. This geometry, adapted from the Qiskit Metal tutorial repository, was selected as a representative benchmark due to its simplicity and familiarity within the superconducting device community. The corresponding eigenmode simulation results are presented in Fig. 2f. In this configuration, the transmon was assigned a Josephson inductance of 11 nH, and the resonator length was 6 mm (excluding the coupling region). In these simulations, we compute the linearized qubit frequency. This parameter plays a key role in the energy participation ratio (EPR) method used to extract circuit Hamiltonian parameters [10], which is discussed further in Appendix B. The solvers again exhibit close agreement as the mesh is refined. At the highest DoF, Palace (MMR, O4) arrived at 6.217 GHz for the transmon and 9.503 GHz for the resonator, while Palace (AMR, O2) arrived at 6.225 GHz and 9.507 GHz, respectively. COMSOL yielded 6.218 GHz (0.032%) and 9.500 GHz (0.112%), and Ansys HFSS produced 6.221 GHz (0.042%) and 9.507 GHz (0.000%) for the qubit and resonator modes, respectively.
@@
The final benchmark design, consisting of two capacitively coupled transmons, is shown in Fig. 2g, with the corresponding eigenmode simulation results presented in Fig. 2h. This configuration, adapted from the Qiskit Metal tutorial repository, was selected because qubit–qubit coupling is a fundamental aspect of circuit quantum electrodynamics (cQED) and quantum computing architectures. In this design, transmon L and transmon R were assigned Josephson inductances of 13 nH and 9 nH, respectively. As in the previous cases, the simulations solve for the linearized transmon frequencies, corresponding to the harmonic eigenmodes of the coupled system. At the highest DoF, the Palace (MMR, O4) solver arrived at 5.747 GHz and 6.902 GHz for transmon L and transmon R, respectively, while Palace (AMR, O2) produced 5.754 GHz and 6.911 GHz. COMSOL yielded 5.745 GHz (0.156%) and 6.900 GHz (0.159%), while Ansys HFSS produced 5.754 GHz (0.000%) and 6.911 GHz (0.000%). Note that we rounded the percentages to three decimal places.
@@
These results confirm that all solvers converge towards nearly identical eigenfrequencies, demonstrating that Palace achieves parity with commercial FEM tools. As shown in Table 1, the total DoF achieved differs significantly between solvers, particularly for the Ansys HFSS simulations. This variation primarily arises from hardware limitations. The Ansys simulations were performed on a local workstation equipped with approximately 760 GB of RAM. During benchmarking, each simulation was allowed to proceed until available memory was exhausted and the solver could no longer store the matrix system after several iterations of mesh refinement. The same procedure was followed for COMSOL, although higher DoF values were reached. This difference likely reflects variations in how each solver internally stores and manages matrix systems, which is information inaccessible to users. In contrast, the Palace simulations were executed on Bunya, the high-performance computing (HPC) cluster at the University of Queensland [24]. The HPC allowed us to utilize larger memory resources, enabling us to simulate with substantially higher DoF. This demonstrates the advantage of HPC-based simulation workflows such as SQDMetal, which can exploit scalable computing resources to achieve improved numerical resolution and accuracy.
@@ III.2 Capacitance Convergence Results
Having established agreement between solvers for eigenmode simulations, we performed electrostatic simulations to verify consistency in the extraction of capacitance values between various solvers. These simulations were conducted for the design of a transmon capacitively coupled to a resonator, shown in Fig. 2e. A detailed view of the transmon and the resonator coupler is presented in Fig. 3a, where the key structural elements are labeled: upper capacitor pad (u), lower capacitor pad (b), readout resonator (r), and ground plane (g). Fig. 3b presents the mutual capacitances extracted from the simulations as required to compute system parameters such as the qubit-resonator couplings and charging energies [25]. These capacitances are also required when using the lumped oscillator model (LOM) [26]. For this convergence study, we compare results from three solvers: the Ansys quasi-static 3D (Q3D) solver using AMR O2, the COMSOL Multiphysics electrostatic solver using MMR O2, and the Palace electrostatic solver using AMR O2. In Fig. 3b, orange, red, and dark blue points represent simulations performed with Ansys, COMSOL, and Palace, respectively, plotted as a function of the notional mesh element size.

### v2

@@ I Introduction
The field of superconducting quantum circuits has emerged as one of the leading theoretical and experimental platforms in the pursuit of fault-tolerant quantum computing [arute_quantum_2019, wu_strong_2021]. Over the past two decades, this maturing field has experienced significant growth, highlighted by the evolving sophistication of experimental techniques and circuit design to improve qubit coherence times [kjaergaard_superconducting_2020]. As the field continues to develop, focus is shifting from understanding the physics underlying superconducting circuit behaviour [nobel] to addressing engineering challenges, such as the design and optimization of device geometries [levenson-falk_review_2024]. Computational tools such as electromagnetic simulation software have become increasingly important for optimizing superconducting circuit design [shammah_open_2024, levenson-falk_review_2024]. Numerous frameworks have been proposed to simulate device behaviour and extract the circuit Hamiltonian using electromagnetic solvers. Among the most widely adopted are the lumped oscillator model (LOM) [minev_circuit_2021, chitta_computer-aided_2022], blackbox quantization (BBQ) [nigg_black-box_2012], and the energy participation ratio (EPR) [minev_energy-participation_2021], which employ electrostatic, driven frequency domain, and eigenmode simulations, respectively. These approaches enable designers to evaluate, refine, and optimize circuit parameters prior to fabrication. However, performing such simulations remains challenging for researchers and engineers with limited access to proprietary electromagnetic solvers such as Ansys HFSS or COMSOL.
@@
  • Qiskit Metal [QiskitMetal] — an open-source framework developed by IBM for the design of superconducting quantum chips and devices.
  • Gmsh [geuzaine_gmsh_2009] — an open-source finite element mesh generator developed by Christophe Geuzaine and Jean-François Remacle for creating 3D meshes and models.
  • Palace (Parallel Large-scale Computational Electromagnetics) [palace] — an open-source electromagnetic solver developed by Amazon Web Services (AWS) for HPC environments, enabling faster and more accurate large-scale simulations.
  • ParaView [paraview] — an open-source visualization and post-processing tool developed by Kitware for analyzing and rendering finite element simulation results.
@@ II Workflow
The operation of the SQDMetal API is shown in Fig. 1, providing a diagrammatic representation of our workflow. The process begins with Qiskit Metal, where the superconducting device is designed. Qiskit Metal features an extensive component library and enables users to programmatically develop circuit layouts in Python via Jupyter notebooks [QiskitMetal]. Qiskit Metal also supports community-driven development, allowing users to create and share custom components that can be integrated into new designs [shanto_squadds_2024].
@@
After meshing, the SQDMetal API transfers the generated mesh file to Palace to perform finite-element electromagnetic simulations. Palace is developed on top of MFEM, an open-source high-performance C++ library for finite-element discretization, and is designed to operate across a wide range of platforms from laptops to supercomputers [anderson_mfem_2021]. While proprietary software packages such as COMSOL Multiphysics and Ansys HFSS also support finite-element electromagnetic simulations on HPCs, there are few open-source alternatives. Palace addresses this gap by providing scalable, open-source computational electromagnetics capabilities. Within SQDMetal, we have integrated Palace’s functionality to perform the following types of simulations:
@@ III Mesh Convergence Study
where the last two designs were taken from the Qiskit Metal tutorial repository. All geometric and material parameters for each device were identical across solvers to ensure a fair comparison. The first two designs were defined on a $500\upmu\text{m}$-thick silicon substrate, while the third and fourth designs, used a $750\upmu\text{m}$-thick substrate. In all cases, the silicon substrate was assigned a relative permittivity of 11.45, a value commonly used for silicon at cryogenic temperatures [krupka_measurements_2006]. The superconducting regions were modeled as zero-thickness surfaces with perfect electric conductor (PEC) boundary conditions. Each chip was enclosed in a rectangular box whose interior volume was defined as vacuum, with all bounding surfaces set to PEC to emulate placement within a metallic sample holder. The enclosing box was twice the height of the chip, while its width and length were 20% larger than the chip. The substrate was placed on the floor in the centre of the enclosing box, thereby grounding the bottom of the device. For the first two designs, the ends of the feedline were terminated with 50 $\Omega$ lumped ports to prevent signal reflections and ensure well-defined boundary conditions.
@@
Decreasing the size of mesh elements allows vector fields (e.g. the electric field in an eigenmode simulation) and scalar fields (e.g., the electric potential in an electrostatic simulation) to be captured with higher spatial resolution, resulting in more accurate solutions. For this benchmark study, we employ exclusively tetrahedral mesh elements. Within each element, the field quantities are interpolated using basis (shape) functions that approximate the field variation across the element [polycarpou_introduction_2006, jin_theory_2010]. Higher-order basis functions improve the accuracy of field interpolation but increase computational cost (in both memory and time). The total number of global unknowns that an electromagnetic solver must compute are referred to as the degrees of freedom (DoF). The DoF directly determine the size of the global matrix system that is constructed at the beginning of the simulation and thus is an indicator of the computational resources required. Specifically, the size of the system matrices for eigenmode or electrostatic simulations is $n\times n$, where $n$ is the DoF. Increasing the number of mesh elements and order of the basis functions both increase the DoF.
@@
When implementing manual mesh refinement (MMR), the user increases the DoF by visually inspecting the electric field distribution and locally decreasing the mesh element size, that is, increasing the mesh element density in regions of strong electric field magnitude. SQDMetal provides specific commands that allow users to selectively refine regions of a device or individual circuit components. However, achieving convergence of a given parameter through MMR can be time-consuming and requires considerable user intuition. The alternative approach, adaptive mesh refinement (AMR), requires less user input while achieving excellent accuracy. In this method, the solver automatically refines the mesh by computing an error indicator that estimates the local deviation between the numerically computed electric field within an element and a smoothed field value calculated in post-processing [zienkiewicz_simple_1987, nicaise_zienkiewiczzhu_2005]. Elements exceeding a user-defined error threshold are subdivided into smaller elements until the estimated error falls below that threshold. AMR requires a reasonably well-refined initial seed mesh to be defined by the user in order to produce the most accurate results [levenson-falk_review_2024]. For both refinement schemes in our study, the polynomial order of the basis functions is held constant; therefore, the convergence behaviour observed in this study arises solely from changes in the mesh element size within regions of high electric field.
@@ III.1 Eigenmode Convergence Results
The results of the eigenmode mesh convergence study for each design are presented in Fig. 2. Each panel in the figure shows a schematic of the device geometry alongside its corresponding plot of resonant frequency as the DoF are increased. In the schematics, blue regions depict superconducting metals, modelled as PEC, while gold regions indicate the exposed silicon substrate. The red rectangles between the capacitor pads are Josephson junctions modelled as lumped element ports. Each point in the plots represents a single eigenmode simulation conducted at a given DoF.
The data points in Fig. 2 correspond to results from different solvers and mesh refinement strategies. Results for Palace simulations using AMR and second order basis functions (AMR, O2) are shown as dark blue circles, while simulations conducted using Palace and MMR with fourth order basis functions (MMR, O4) are shown as light blue circles. Ansys HFSS simulations using AMR and second order basis functions (AMR, O2) are plotted in orange, while COMSOL Multiphysics simulations using MMR and second order basis functions (MMR, O2) are in red. For this study, we define a notional mesh element size as $m=\text{DoF}^{-1/3}$ [palace_tutorial] where $m\to 0$ as $\text{DoF}\to\infty$ to approach the continuum limit. The simulated resonant frequencies are plotted against this quantity to illustrate convergence of each solver toward a common frequency value. The corresponding total DoF for each simulation is indicated along the top axis of each plot for reference.
@@
The first benchmark design, shown in Fig. 2a, is a single coplanar waveguide (CPW) resonator. This geometry was selected as a baseline case, since a $\lambda/2$ resonator is a common component in superconducting circuits. The corresponding plot in Fig. 2b shows the convergence of the simulated resonant frequency. The resonator was designed to operate at 10 GHz based on the analytical expression for a half-wavelength CPW resonator [goppl_coplanar_2008]:
@@
where $f$ is the resonant frequency of the fundamental mode, $c$ is the speed of light in vacuum, $\ell$ is the length of the resonator and $\varepsilon_{\text{eff}}$ is the effective permittivity that accounts for field penetration into both vacuum and the substrate. The effective permittivity was approximated as $\varepsilon_{\text{eff}}=(\varepsilon_{\text{Si}}+\varepsilon_{\text{vac}})/2$, where $\varepsilon_{\text{Si}}=11.45$ [krupka_measurements_2006] and $\varepsilon_{\text{vac}}=1$. Substituting $f=10$ GHz yields a resonator length of 6.012 mm.
@@
The second benchmark design, a coplanar waveguide (CPW) resonator capacitively coupled to a feedline, is shown in Fig. 2c, with corresponding eigenmode simulation results presented in Fig. 2d. This configuration introduces an additional degree of complexity relative to the single-resonator case via a capacitive coupling to a transmission feedline. For this $\lambda$/2 resonator, a length of 8.475 mm was used, which, according to Eq. 1, yields an expected fundamental frequency of approximately 7.089 GHz. As shown in Fig. 2d, at the highest DoF, Palace (MMR, O4) arrived at 7.095 GHz, and Palace (AMR, O2) produced 7.094 GHz. COMSOL Multiphysics and Ansys HFSS produced frequencies of 7.092 GHz (0.028%) and 7.101 GHz(0.097%), respectively.
@@
The next benchmark design, shown in Fig. 2e, consists of a transmon capacitively coupled to a coplanar waveguide (CPW) resonator. This geometry, adapted from the Qiskit Metal tutorial repository, was selected as a representative benchmark due to its simplicity and familiarity within the superconducting device community. The corresponding eigenmode simulation results are presented in Fig. 2f. In this configuration, the transmon was assigned a Josephson inductance of 11 nH, and the resonator length was 6 mm (excluding the coupling region). In these simulations, we compute the linearized qubit frequency. This parameter plays a key role in the energy participation ratio (EPR) method used to extract circuit Hamiltonian parameters [minev_energy-participation_2021], which is discussed further in Appendix B. The solvers again exhibit close agreement as the mesh is refined. At the highest DoF, Palace (MMR, O4) arrived at 6.217 GHz for the transmon and 9.503 GHz for the resonator, while Palace (AMR, O2) arrived at 6.225 GHz and 9.507 GHz, respectively. COMSOL yielded 6.218 GHz (0.112%) and 9.500 GHz (0.074%), and Ansys HFSS produced 6.221 GHz (0.064%) and 9.507 GHz (0.000%) for the qubit and resonator modes, respectively.
@@
The final benchmark design, consisting of two capacitively coupled transmons, is shown in Fig. 2g, with the corresponding eigenmode simulation results presented in Fig. 2h. This configuration, adapted from the Qiskit Metal tutorial repository, was selected because qubit–qubit coupling is a fundamental aspect of circuit quantum electrodynamics (cQED) and quantum computing architectures. In this design, transmon L and transmon R were assigned Josephson inductances of 13 nH and 9 nH, respectively. As in the previous cases, the simulations solve for the linearized transmon frequencies, corresponding to the eigenmodes of the coupled system. At the highest DoF, the Palace (MMR, O4) solver arrived at 5.747 GHz and 6.902 GHz for transmon L and transmon R, respectively, while Palace (AMR, O2) produced 5.754 GHz and 6.911 GHz. COMSOL yielded 5.745 GHz (0.156%) and 6.900 GHz (0.159%), while Ansys HFSS produced 5.754 GHz (0.000%) and 6.911 GHz (0.000%). Note that we rounded the percentages to three decimal places.
@@
These results confirm that all solvers converge towards nearly identical eigenfrequencies, demonstrating that Palace achieves parity with commercial FEM tools. As shown in Table 1, the total DoF achieved differs significantly between solvers, particularly for the Ansys HFSS simulations. This variation primarily arises from hardware limitations. The Ansys simulations were performed on a local workstation equipped with approximately 768 GB of RAM. See Appendix D for details on the computers/hardware used in this study. During benchmarking, each simulation was allowed to proceed until available memory was exhausted and the solver could no longer store the matrix system after several iterations of mesh refinement. The same procedure was followed for COMSOL, although higher DoF values were reached. This difference likely reflects variations in how each solver internally stores and manages matrix systems, which is information inaccessible to users. In contrast, the Palace simulations were executed on Bunya, the high-performance computing (HPC) cluster at the University of Queensland [bunya]. Using the HPC allowed us to utilize larger memory resources, enabling us to simulate with substantially higher DoF. This demonstrates the advantage of HPC-based simulation workflows such as SQDMetal, which can exploit scalable computing resources to achieve improved numerical resolution and accuracy.
@@ III.2 Capacitance Convergence Results
Having established agreement between solvers for eigenmode simulations, we performed electrostatic simulations to verify consistency in the extraction of capacitance values between various solvers. These simulations were conducted for the design of a transmon capacitively coupled to a resonator, shown in Fig. 2e. A detailed view of the transmon and the resonator coupler is presented in Fig. 3a, where the key structural elements are labeled: upper capacitor pad (u), lower capacitor pad (b), readout resonator (r), and ground plane (g). Fig. 3b presents the mutual capacitances extracted from the simulations as required to compute system parameters such as the qubit-resonator couplings and charging energies [blais_circuit_2021]. These capacitances are also required when using the lumped oscillator model (LOM) [groszkowski_scqubits_2021]. For this convergence study, we compare results from three solvers: the Ansys quasi-static 3D (Q3D) solver using AMR O2, the COMSOL Multiphysics electrostatic solver using MMR O2, and the Palace electrostatic solver using AMR O2. In Fig. 3b, orange, red, and dark blue points represent simulations performed with Ansys, COMSOL, and Palace, respectively, plotted as a function of the notional mesh element size.